I am require $11,000 in two years to return to college to get an MBA degree. How much money should I ask my father for now so if I invests it at 9% interest compounded continuously, I will have enough for school? Round answer to the nearest dollar amount.

Answer by stanbon(75887)

I am require $11,000 in two years to return to college to get an MBA degree. How much money should I ask my father for now so if I invests it at 9% interest compounded continuously, I will have enough for school? Round answer to the nearest dollar amount.
---------
A(t)=Pe^(rt)
11000 = Pe^(0.09*2)
P = 11000/e^(0.18)
P = 9187.97
Rounding: P =$9188
-----------
Cheers,
Stan H.